OpenStudy (anonymous):

What is the solution to the equation 256 = (1/2x - 3)2?

OpenStudy (pfenn1):

Is this your equation?\[256 = (\frac12x - 3)^2\]

OpenStudy (anonymous):

yea

OpenStudy (pfenn1):

Take the square root of each side.

OpenStudy (anonymous):

so take the square root of 256 and (1/2x-3)^3

OpenStudy (pfenn1):

Yes, but I thought the right side was (1/2x-3)^2?

OpenStudy (anonymous):

it is, that was a mistake

OpenStudy (anonymous):

(1/2x-3)^2 is right

OpenStudy (pfenn1):

Is it\[ (\frac12x-3)^2 \text{ or }(\frac{1}{2x}-3)^2\]

OpenStudy (pfenn1):

Well, anyway, what is \[\sqrt{256}=\sqrt{(\frac12x-3)^2}\]

OpenStudy (precal):



OpenStudy (precal):